Solve the following; a) If the cost for 6 oranges is 60 rupees, what will be the cost of 6 ¾ dozen oranges? b) How many miles are there between two cities, if the distance is represented by a 4.2 inch line on a map having a scale of 1 inch to 38 miles? c) |3𝑥 − 5| = |−7 − 8𝑥| d) 𝑥 2 + 4𝑥 − 21 , using factoring e) Find the midpoint and the distance of the line segment connecting the points (-3,-7 and 5, -22)


1
Expert's answer
2021-04-13T17:34:39-0400

a)

cost of 6 oranges is 60 rupees

cost of 1 orange is "\frac{60}{6}=10" rupees

1 dozen=12, therefore the cost of "6\frac{3}{4}" dozen oranges is

"10⋅6\frac{3}{4}⋅12=10⋅\frac{27}{4}⋅12=10⋅27⋅3=810" rupees

b)

"4.2\u22c538=159.60" miles

c)

"|3x-5|=|-7-8x|"

1) "3x-5=-7-8x"

"3x+8x=5-7"

"11x=-2"

"x=-\frac{2}{11}"

2) "3x-5=-(-7-8x)"

"3x-5=7+8x"

"3x-8x=5+7"

"-5x=12"

"x=-\frac{12}{5}"

3) "-(3x-5)=-7-8x"

"-3x+5=-7-8x"

"-3x+8x=-5-7"

"5x=-12"

"x=-\frac{12}{5}"

4) "-(3x-5)=-(-7-8x)"

"3x+5=7+8x"

"-3x-8x=-5+7"

"-11x=2"

"x=-\frac{2}{11}"

d)

"x^2+4x-21=0"

"x^2+7x-3x-21=0"

"x(x+7)-3(x+7)=0"

"(x-3)(x+7)=0"

"x_1=3" and "x_2=-7"

e)

midpoint is:

"(\frac{(x_2+x_1)}{2}, \frac{(y_2+y_1)}{2})=(\frac{5-3}{2}, \frac{-22-7}{2})=(1, -14.5)"